OpenStudy (anonymous):

place the answer in scentific notation, when appropriate and simplify the units. (1.33)sin35.0=(1.50)sinθ 35.0 is in degrees

OpenStudy (fifciol):

(1.33)sin35.0/(1.50)=sin theta

OpenStudy (fifciol):

sin 35 =0.573576436

OpenStudy (fifciol):

sin theta= 0.573576436*1.33/1.5=0.508571106 theta=30.57 degrees

OpenStudy (anonymous):

thanks i got it

OpenStudy (fifciol):

yw:)

OpenStudy (theeric):

And the answer had only three significant figures, right? I'm just saying this in case someone looks through closed questions or finds it through Google. So, if the calculations were right, the answer should be \(30.6^\circ\) with proper significant figures. For significant figures, all you need to know is 1. how to count them, and 2. how you choose how many after addition/subtraction or multiplication/division.

OpenStudy (anonymous):

what r significant figures

OpenStudy (anonymous):

do u know rosie